Given that log2=0.3010, log7=0.845 and log5=0.699
Evaluate log45

Answer by ikleyn(53749)

.
Given that log2=0.3010, log7=0.845 and log5=0.699
Evaluate log45
~~~~~~~~~~~~~~~~~~~~~


45 = 9*5 = 3*3*5.


THEREFORE,   log(45) = log(3*3*5) = log(3) + log(3) + log(5) = 2*log(3) + log(5).


In this problem,  given  log(7) = 0.845  is  IRRELEVANT.


Log(3)  should be given in order for the solution be possible,  but it is  NOT  GIVEN,   and,   therefore,
the solution  CAN  NOT  be obtained.

-------------------

As posed,  given problem is  a   F A K E.